OpenStudy (anonymous):

plzzzz help: find all values of theta in the interval 0

OpenStudy (cwrw238):

3cos2x + 2sinx + 1 = 0 use the relation cos2x = 1 - 2 sin^2x to get an equation is sinx

OpenStudy (anonymous):

wait is tht an identity? cos2x = 1 - 2 sin^2x

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

so replace cos3x by 1 - 2sin^2 x and you'll get a quadratic equation in sin x

OpenStudy (cwrw238):

* replace cos 2x

OpenStudy (anonymous):

so um cos2x=3cos2theta?

OpenStudy (cwrw238):

oh yes i was lazy so i put x instead of theta

OpenStudy (anonymous):

what do i do wit the rest? 2sin theta+1=0?

OpenStudy (cwrw238):

ok ill write everything in theta cos 2 theta = 1 - 2sin^2 theta so 3(1 - 2sin^2 theta) + 2 sin theta + 1 = 0 now simplify this to get a quadratic in sin theta

OpenStudy (anonymous):

i got sin^2theta=1

OpenStudy (anonymous):

correct so faR?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

i got theata=90?

OpenStudy (cwrw238):

hold on

OpenStudy (anonymous):

okk

OpenStudy (cwrw238):

3(1 - 2sin^2 theta) + 2 sin theta + 1 = 0 3 - 6sin^2 theta + 2 sin theta + 1 = 0 6sin^2 theta - 2 sin theta - 4 = 0 3 sin^2 theta - sin theta - 2 = 0

OpenStudy (cwrw238):

did you get that?

OpenStudy (cwrw238):

factoring (3sin theta + 2)(sin theta - 1) = 0

OpenStudy (anonymous):

oh isee wht i did. one sec plz

OpenStudy (anonymous):

y -4? its shuld be positive 3+1=4

OpenStudy (anonymous):

and y r u halfing the 6sin^2theta nd everything else?

OpenStudy (anonymous):

so -3 on both sides udid?

OpenStudy (anonymous):

ok so now i have 3-6sin^2theta+2sin theta+1=0 ---->do i subtract 3 on both sides?

OpenStudy (anonymous):

need the slope

OpenStudy (anonymous):

wait 1-3 is -2

OpenStudy (anonymous):

how did u gt -4?

OpenStudy (cwrw238):

-1 -3 = -4

OpenStudy (anonymous):

wrong:3-6sin^2theta+2sin theta+1=0 is this right:3-6sin^2theta+2sin theta-1=0 i had +1 insteads of -1

OpenStudy (anonymous):

howcom ur doin tht? i nevr did be4

OpenStudy (cwrw238):

if you like you can look at it this way I multiplied each term by -1 in order that first term (the 6 sin^2 theta becomes positive the equation still hodls and is easier to factor

OpenStudy (anonymous):

but my way wont work?

OpenStudy (cwrw238):

what was your way?

OpenStudy (anonymous):